tan(cos^-1 (1/3)) = ??
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

tan(cos^-1 (1/3)) = ??
-------------
Find a terminal point.
cos = x/r = 1/3
y = sqrt(r^2 - x^2) = sqrt(8)
Terminal point is (1,sqrt(8))
-----
tan = y/x
= sqrt(8)

Answer by stanbon(75887)

tan(cos^-1 (1/3)) = ??
--------
Since the cos is 1/3, x = 1 and r = 3,
Then y = sqrt[3^2-1^2] = sqrt(8)
----
Ans: tan = y/x = sqrt(8)
--------------
Cheers,
Stan H.
------------